# Skykatana: a scalable framework to construct sky masks for the Vera Rubin Observatory and large astronomical surveys

### v1

@@ 1 Introduction
The rapid growth of modern wide-field imaging surveys has fundamentally changed the way astronomers characterize the observable Universe. Projects such as the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST), the Subaru Hyper Suprime-Cam (HSC) survey, and the Dark Energy Survey (DES) have achieved unprecedented sky coverage and depth, producing petabyte-scale datasets that must be rigorously characterized in terms of angular completeness and data quality. For most scientific applications, ranging from galaxy clustering and cosmic shear measurements to stellar population and transient studies, it is essential to know whether the survey data at every point on the sky meet the required quality criteria. This information is encoded in spatial masks that exclude regions contaminated by bright stars, diffraction spikes, detector artifacts, poor seeing, astrophysical foregrounds, or incomplete coverage, and forms the backbone of most selection functions in modern surveys.
Traditional geometric mask frameworks such as MANGLE (hamilton2004, swanson2008) represent the sky as unions of spherical polygons with analytic boundaries, achieving exact area computation, but requiring complex data structures, tolerance to numerical instability, and expensive spherical operations that scale poorly with the number of polygons. Pixelization-based approaches, most notably HEALPix (gorski2005), offer a simpler and faster alternative by discretizing the sphere into pixels of equal area. HEALPix is ideally suited for dense, full-sky scalar fields (e.g., maps of temperature, intensity, or depth), but becomes inefficient for sparse boolean maps because memory use scales linearly with the total number of pixels. The HEALSparse library (hsppypi, Dark Energy Science Collaboration) addressed this limitation by introducing a hierarchical ”coverage/sparse” representation, where only the pixels actually present in the mask are stored, preserving exact geometry while reducing memory and I/O costs by orders of magnitude.
The Multi-Order Coverage framework (MOC; fernique2015) allows compact, hierarchical representation of arbitrary sky regions, and is widely used visualizing survey footprints, catalog coverage, and boolean region operations. The MOCpy Python library (boch2019) implements an interface for manipulating MOCs in pure Python. However, MOCpy is primarily focused on region-level operations and is not optimized for high-resolution boolean masking of extended exclusion zones (e.g., around bright stars) across billions of pixels.
@@
Building upon this foundation, Skykatana provides a general-purpose modular framework to construct, combine, and visualize such sparse boolean maps at survey scale. It introduces the concept of mask pipelines, composed of independent stages (e.g. footprints, bright-star masks, depth thresholds, or extended-source masks), each represented internally as a HEALSparse map. These stages can be generated from catalogs, property maps, or analytic geometric primitives, and then logically combined through union, intersection, or subtraction into a final mask suitable for scientific use. Skykatana emphasizes both computational efficiency and workflow clarity, enabling users to generate reproducible, documented masking procedures that scale from small fields to the entire Rubin footprint. Skykatana is designed to operate efficiently in distributed environments and with HATS/LSDB schemas (caplar2025), where astronomical catalogs are hierarchically partitioned in HEALPix cells. The code implements pixel streaming algorithms that enable processing large areas of the sky, automatically query large datasets that would be costly to search locally, and stream the resulting pixel sets directly into bit-packed maps. These features enable the generation of high-resolution masks of the entire Rubin footprint directly in the Rubin Science Platform (RSP) with modest computational resources.
@@
In this paper, we present the design and capabilities of Skykatana, describe its internal algorithms and data model, and demonstrate two end-to-end applications: (i) a Subaru HSC–WISE composite mask, combining footprint, depth, and bright-star stages; (ii) a bright-star mask for the Rubin Observatory over the Wide-Fast-Deep survey area, constructed on demand via HATS/LSDB queries using empirical magnitude–radius relations derived from Rubin Data Preview 1 (DP1) data.
@@ 2.1.1 Implementation and distribution
Skykatana is implemented entirely in Python ($\geq~3.10$) and distributed as an installable package on PyPI under the name skykatana. It can be installed with a single command (pip install skykatana) and requires the HATS/LSDB libraries along with standard scientific dependencies (numpy, pandas, astropy, healsparse, and mocpy). The package is fully open source under the MIT license, hosted on GitHub, and accompanied by comprehensive online documentation and example notebooks.
@@ 2.2.3 Catalog-driven builders
In some cases, pixelizing extensive surveys can be impractical due to the sheer size of the catalogs, or technical difficulties to estimate precise coverage regions. In some cases like Rubin the dataset will be too large to fit into memory without a significant computational infrastructure. To accommodate such cases, the code implements a method (build_foot_mask()) to directly pixelize a dense catalog of discrete sources. When a dataset has been partitioned with the HATS/LSDB framework it is already organized into roughly equal-size partitions (or HATS pixels) and materialized as a set of parquet files which can be loaded and processed in turn by one or more workers of a Dask cluster. Skykatana wraps its pixelization logic as a function that is mapped to every partition, sent for execution across the cluster, and collects (reduces) the pixels at the end. In this way, an enormous task is divided into smaller pieces, each distributed across partitions, achieving out-of-core scalability with limited resources.
@@ 2.3 Star mask builder: scaling pixelization on demand
We also include a detailed description of a special method to construct bright star masks. build_star_mask_online() is designed to build an exclusion mask by querying an external stellar catalogs (e.g. Gaia in HATS/LSDB version) on demand over an arbitrary sky region, then pixelizing circular exclusion zones around each star using a given radius relation (for instance, the empirical radius–magnitude relation from Sec. 3). Compared to precomputed, monolithic stellar masks, this online approach offers several advantages: (i) users can tailor magnitude limits, radius prescriptions, and sky coverage without regenerating global products; (ii) throughput and memory are bounded by chunked processing of the target region; (iii) avoids downloading and querying large catalogs locally, which can involve significant amount of time and resources; (iv) the method naturally supports distributed execution (e.g. within the Rubin Science Platform, see LDM542-RSPDesign, LSE319-RSPVision, omullane2024).
@@ 2.3.1 Search-stage definition and MOC construction
Users provide a search stage that defines where stars should be queried and ultimately masked out. Most often, this is the survey footprint or a quality-cut mask. The stage is converted into its multi-order coverage (MOC) representation at a user-specified maximum order $o_{\max}$, yielding a set of nested HEALPix cells $\mathcal{M}_{\mathrm{search}}=\{(o_{i},p_{i})\}$ whose union approximates the stage to subpixel precision.
@@ 2.4.1 MOC and static quick-look views
Each stage can be visualized through its corresponding Multi-Order Coverage (MOC) representation (using the plot_moc() method). The code computes the minimal MOC footprint at the desired maximum order that is worth plotting, given the chosen viewport, and displays it as a set of filled HEALPix cells in a WCS projection. Not less important, it also clips pixels to the current viewport, allowing fast rendering of large footprints. It also supports multiple overlaid stages, enabling direct comparison between. These plots provide a rapid diagnostic of the stage alignment, coverage, and relative contribution of each component to the final mask.
@@ 2.4.2 Visualization by random points
A complementary mask visualization method is through random point sampling. Skykatana can uniformly draw random coordinates within the valid (unmasked) region of a stage or of the final combined mask. These randoms are then plotted to verify uniformity and continuity, and to identify potential discontinuities or holes introduced by masking operations. Because only the random points are displayed rather than the underlying boolean map, this method allows visual inspection of very high-order masks with negligible memory cost and high performance.
@@ 2.4.3 Interactive exploration
For interactive inspection, Skykatana integrates seamlessly with the ipyaladin Jupyter widget. Stage masks can be automatically exported as MOCs and displayed in an ipyaladin viewer layer, enabling users to pan, zoom, and overlay catalogs or images directly from the notebook. This interface is particularly useful for examining the correctness of star-exclusion radii, verifying geometry boundaries, or cross-checking against sky images.
@@ 2.6 IO and Interoperability
In addition to map stages, the saved directory also contains a JSON file which serves two purposes : (i) record the metadata produced by each of the ’build’ method calls used to generate the corresponding stages, including also the number of valid pixels and total area; and (ii) hold information to preserve state and reattach the stages under their original names when read back.
@@ 3.1 Data pre-processing
To calibrate the stellar exclusion radii, we used Rubin DP1 photometry (RTN095-DP1) within the ECDFS. Our first goal was to construct a background distribution of sources with approximately uniform spatial completeness. For this, we made use of the deepCoadd_psf_maglim_consolidated_map_weighted_mean property maps available in the Rubin Science Platform, which store the depth at HEALPix resolution as HEALSparse maps. These property maps allow us to determine the maximum depth reached within the ECDFS for each band, and we adopted this value as the faintest usable magnitude for the DP1 sources. These limits correspond to 25.39, 27.38, 27.12, 26.53, 25.67, and 23.77 in the u, g, r, i, z, y bands, respectively. Next, to remove shallow border regions and ensure spatial uniformity, we used the build_prop_mask function of Skykatana to generate a boolean mask selecting only those pixels where the local depth exceeds a band-dependent threshold. The thresholds were 24.5, 26.5, 26.0, 25.5, 25.0, and 23.0 in u, g, r, i, z, and y, respectively. Only DP1 sources within these masks were included in the analysis.
@@
A well-known complication in LSST images is that bright saturated stars produce numerous artificial detections through deblender-induced fragmentation of both the stellar core and the scattered-light halo. These spurious detections produce pronounced overdensities around bright stars. The LSST blendedness_flag is raised when the pipeline is unable to compute a reliable blendedness metric for a given detection. This typically occurs under non-ideal photometric or geometric conditions such as saturated stellar cores, extremely steep flux gradients, irregular footprints, or severe contamination where the algorithm cannot robustly determine the relation between the flux of the source and that of its inferred parent source. In practice, this flag correlates strongly with spurious deblender fragments around bright stars, and selecting only objects with blendedness_flag == False effectively removes the majority of these artifacts. Although this cut eliminates some real faint galaxies projected over bright halos, its impact on the averaged profiles is modest. For this reason, we carried out two parallel analyses: one including all DP1 sources and thus preserving the halo-induced overdensities, and one excluding halo artifacts via blendedness_flag == False. Figure 2 shows two representative mean profiles. When all sources are included, the profile exhibits a spurious central excess, followed by a depletion where the star suppresses faint detections, and eventually a secondary excess associated with halo fragmentation. When halo artifacts are removed, the profile increases smoothly and monotonically toward unity. These qualitative differences motivate the distinct definitions of stellar radius described below.
Figure 3: Fitted radius–magnitude relations for all LSST bands in the halo and no-halo cases. For the halo case, three power-law segments are fitted per band, with boundaries at $G_{\mathrm{Gaia}}=13.6$ and $G_{\mathrm{Gaia}}=15.0$, capturing the transition from strong halo contamination to the weak-halo regime. The no-halo case is well described by a single power law per band. Best-fit parameters are listed in Tables 1 and 2.

### v2

@@ 1 Introduction
The rapid growth of modern wide-field imaging surveys has fundamentally changed the way astronomers characterize the observable Universe. Projects such as the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezić et al. 2019, Rubin Observatory Science Pipelines Developers 2025), the Subaru Hyper Suprime-Cam (HSC; Aihara et al. 2018) survey, and the Dark Energy Survey (DES; Porredon et al. 2022a, Porredon et al. 2022b, Rodríguez-Monroy et al. 2022, Rodríguez-Monroy et al. 2025) have achieved unprecedented sky coverage and depth, producing petabyte-scale datasets that must be rigorously characterized in terms of angular completeness and data quality. For most scientific applications, ranging from galaxy clustering and cosmic shear measurements to stellar population and transient studies, it is essential to know whether the survey data at every point on the sky meet the required quality criteria. This information is encoded in spatial masks that exclude regions contaminated by bright stars, diffraction spikes, detector artifacts, poor seeing, astrophysical foregrounds, or incomplete coverage, and forms the backbone of most selection functions in modern surveys.
Traditional geometric mask frameworks such as MANGLE (Hamilton and Tegmark (2004), Swanson et al. (2008)) represent the sky as unions of spherical polygons with analytic boundaries, achieving exact area computation, but requiring complex data structures, tolerance to numerical instability, and expensive spherical operations that scale poorly with the number of polygons. Pixelization-based approaches, most notably HEALPix (Górski et al., 2005), offer a simpler and faster alternative by discretizing the sphere into pixels of equal area. HEALPix is ideally suited for dense, full-sky scalar fields (e.g., maps of temperature, intensity, or depth), but becomes inefficient for sparse boolean maps because memory use scales linearly with the total number of pixels. The HEALSparse library (Rykoff and Sánchez 2023, Dark Energy Science Collaboration) addressed this limitation by introducing a hierarchical ”coverage/sparse” representation, where only the pixels actually present in the mask are stored, preserving exact geometry while reducing memory and I/O costs by orders of magnitude.
The Multi-Order Coverage framework (MOC; Fernique et al. 2015) allows compact, hierarchical representation of arbitrary sky regions, and is widely used visualizing survey footprints, catalog coverage, and boolean region operations. The MOCpy Python library (Boch, 2019) implements an interface for manipulating MOCs in pure Python. However, MOCpy is primarily focused on region-level operations and is not optimized for high-resolution boolean masking of extended exclusion zones (e.g., around bright stars) across billions of pixels.
@@
Therefore, new tools are needed to manage and combine the multiple property maps required to define the angular selection function of future surveys like the NSF-DOE Vera C. Rubin Observatory’s Legacy survey of Space and Time (LSST); Ivezić et al. (2019).
Building upon this foundation, Skykatana provides a general-purpose modular framework to construct, combine, and visualize such sparse boolean maps at survey scale. It introduces the concept of mask pipelines, composed of independent stages (e.g. footprints, bright-star masks, depth thresholds, or extended-source masks), each represented internally as a HEALSparse map. These stages can be generated from catalogs, property maps, or analytic geometric primitives, and then logically combined through union, intersection, or subtraction into a final mask suitable for scientific use. Skykatana emphasizes both computational efficiency and workflow clarity, enabling users to generate reproducible, documented masking procedures that scale from small fields to the entire Rubin footprint. Skykatana is designed to operate efficiently in distributed environments and with HATS/LSDB111docs.lsdb.io/en/latest/ schemas (Caplar et al. 2025), where astronomical catalogs are hierarchically partitioned in HEALPix cells. The code implements pixel streaming algorithms that enable processing large areas of the sky, automatically query large datasets that would be costly to search locally, and stream the resulting pixel sets directly into bit-packed maps. These features enable the generation of high-resolution masks of the entire Rubin footprint directly in the Rubin Science Platform (RSP) instance hosted at https://data.lsst.cloud (O’Mullane et al., 2024) with modest computational resources.
@@
In this paper, we present the design and capabilities of Skykatana, describe its internal algorithms and data model, and demonstrate two end-to-end applications: (i) a Subaru Hyper Suprime-Cam and Widefield Infrared Survey Explorer (HSC–WISE) composite mask, combining footprint, depth, and bright-star stages; (ii) a bright-star mask for the Rubin Observatory over the Wide-Fast-Deep survey area, constructed on demand via HATS/LSDB queries using empirical magnitude–radius relations derived from Rubin Data Preview 1 (DP1; NSF-DOE Vera C. Rubin Observatory 2025b, Malanchev et al. 2025) data.
The primary objectives of this work are:
  1. 1.
    to introduce Skykatana as a scalable framework for constructing and combining sky masks at the resolution and data volumes required by next-generation surveys;
  2. 2.
    to enable on-demand, catalog-driven mask generation through distributed query engines;
  3. 3.
    to provide a reproducible and composable masking logic that can be stored, combined, and reused across analyses;
  4. 4.
    to demonstrate these capabilities using representative masking applications drawn from existing survey data.
Existing masking approaches typically fall into static, survey-specific polygon masks that are difficult to scale or adapt. Skykatana is designed to fill this gap by enabling high-resolution, dynamically generated masks that remain computationally tractable and reproducible at Rubin-scale data volumes. This gap becomes critical in modern survey workflows, where analyses are increasingly executed in cloud-based notebook environments with constrained memory, require distributed processing over billions of pixels, and demand strict reproducibility of mask definitions across teams and analyses. These practical constraints strongly motivate the design choices underlying Skykatana. The demonstration applications presented in this paper based on HSC-WISE masking and Rubin bright-star exclusion are intended to illustrate the capabilities of the Skykatana framework, rather than to provide definitive or universal scientific calibrations.
@@
Design highlights
The Skykatana framework is designed around the following key principles:
  • 1.
    Scalable sparse masking: Angular masks containing billions of HEALPix pixels are represented using sparse, bit-packed encodings and streaming I/O, enabling their use in memory-constrained notebook and cloud environments.
  • 2.
    On-demand, catalog-driven construction: Mask stages are generated directly from distributed catalogs via LSDB/HATS queries, avoiding the need for precomputed full-sky products.
  • 3.
    Chunked, hierarchical workflow: A breadth-first chunking strategy ensures bounded memory usage while scaling naturally with survey area.
  • 4.
    Composable pipeline abstraction: Mask stages can be combined, persisted, and reloaded reproducibly using a unified pipeline model.
  • 5.
    Interoperability and visualization: Native support for MOC export and interactive notebook visualization enables rapid inspection and validation of complex masks.
@@ 2.1.1 Implementation and distribution
Skykatana is implemented entirely in Python ($\geq~3.10$) and distributed as an installable package on PyPI under the name skykatana. It can be installed with a single command (pip install skykatana) and requires the HATS/LSDB libraries along with standard scientific dependencies (numpy, pandas, astropy, healsparse, and mocpy). The package is fully open source under the MIT license, hosted on GitHub, and accompanied by comprehensive online documentation and example notebooks. To ensure reproducibility, the exact version used in this work has been archived with a persistent DOI through Zenodo (DOI: 10.5281/zenodo.18407348). The archive also includes a conda environment file and pinned dependency versions.
@@ 2.2.3 Catalog-driven builders
In some cases, pixelizing extensive surveys can be impractical due to the sheer size of the catalogs, or technical difficulties to estimate precise coverage regions. In some cases like Rubin the dataset will be too large to fit into memory without a significant computational infrastructure. To accommodate such cases, the code implements a method (build_foot_mask()) to directly pixelize a dense catalog of discrete sources. When a dataset has been partitioned with the HATS/LSDB framework it is already organized into roughly equal-size partitions (or HATS pixels) and materialized as a set of parquet files which can be loaded and processed in turn by one or more workers of a Dask cluster (Dask Development Team, 2016). Skykatana wraps its pixelization logic as a function that is mapped to every partition, sent for execution across the cluster, and collects (reduces) the pixels at the end. In this way, an enormous task is divided into smaller pieces, each distributed across partitions, achieving out-of-core scalability with limited resources.
@@ 2.3 Star mask builder: scaling pixelization on demand
We also include a detailed description of a special method to construct bright star masks. build_star_mask_online() is designed to build an exclusion mask by querying an external stellar catalogs (e.g. Gaia in HATS/LSDB version) on demand over an arbitrary sky region, then pixelizing circular exclusion zones around each star using a given radius relation (for instance, the empirical radius–magnitude relation from Sec. 3). Compared to precomputed, monolithic stellar masks, this online approach offers several advantages: (i) users can tailor magnitude limits, radius prescriptions, and sky coverage without regenerating global products; (ii) throughput and memory are bounded by chunked processing of the target region; (iii) avoids downloading and querying large catalogs locally, which can involve significant amount of time and resources; (iv) the method naturally supports distributed execution (e.g. within the Rubin Science Platform, see Dubois-Felsmann et al. 2019, Jurić et al. 2019, O’Mullane et al. 2024).
@@ 2.3.1 Search-stage definition and MOC construction
Users provide a search stage that defines where stars should be queried and ultimately masked out. Most often, this is the survey footprint or a quality-cut mask. The stage is converted into its MOC representation at a user-specified maximum order $o_{\max}$, yielding a set of nested HEALPix cells $\mathcal{M}_{\mathrm{search}}=\{(o_{i},p_{i})\}$ whose union approximates the stage to subpixel precision.
@@ 2.3.4 Milky Way masking during search
2.4 Random point generation
Skykatana generates random points uniformly over masks using a two-pass streaming algorithm that avoids materializing the full list of valid pixels. The mask is iterated in chunks grouped by coverage pixel. In a first pass, the number of valid sparse pixels per chunk is counted, and the requested total number of random points is distributed across chunks via a multinomial draw, proportional to the valid area. In a second pass, parent sparse pixels are sampled uniformly within each chunk and refined to a finer grid (by default nside_randoms=23) by appending random sub-pixel bits, yielding uniform sampling within each sparse pixel. This design ensures bounded memory usage and scalability to high-resolution, highly perforated masks.
We quantify uniformity using a Kolmogorov–Smirnov (K–S) test at the native sparse resolution. Each random point is mapped to its sparse pixel index and converted to a continuous rank statistic over valid pixels. We choose a highly perforated mask (the join of LMC r–band masks including the halo component and faint Gaia stars, see Appendix A). Using $10^{7}$ random points and 20 independent seeds, we obtain K–S statistics $D\sim 2-4\times 10^{-4}$, consistent with the expected $\frac{1}{\sqrt{n}}$ behavior and showing no statistically significant deviation from uniformity.
As an independent test, we compute the angular two-point correlation function $w(\theta)$ using the Landy–Szalay estimator, with independent random catalogs ($N_{D}=2\times 10^{6}$, $N_{R}=2\times 10^{7}$). Over the fitted angular range from $1^{\prime\prime}$ to $1^{\circ}$, $w(\theta)$ is consistent with zero, with $\chi^{2}/\text{dof}=0.16~(p=0.99998)$ for the null hypothesis $w(\theta)=0$.
@@ 2.5.1 MOC and static quick-look views
Each stage can be visualized through its MOC representation (using the plot_moc() method). The code computes the minimal MOC footprint at the desired maximum order that is worth plotting, given the chosen viewport, and displays it as a set of filled HEALPix cells in a World Coordinate System (WCS) projection. Not less important, it also clips pixels to the current viewport, allowing fast rendering of large footprints. It also supports multiple overlaid stages, enabling direct comparison between. These plots provide a rapid diagnostic of the stage alignment, coverage, and relative contribution of each component to the final mask.
@@ 2.5.2 Visualization by random points
A complementary mask visualization method is through random point sampling. As said before, Skykatana can uniformly draw random coordinates within the valid (unmasked) region of a stage or of the final combined mask. These randoms are then plotted to verify uniformity and continuity, and to identify potential discontinuities or holes introduced by masking operations. Because only the random points are displayed rather than the underlying boolean map, this method allows visual inspection of very high-order masks with negligible memory cost and high performance.
@@ 2.5.3 Interactive exploration
Skykatana integrates seamlessly with the ipyaladin Jupyter widget, a Python interface to the Aladin Lite interactive sky atlas (Boch and Fernique 2014), enabling interactive visualization and inspection of mask stages directly within Jupyter notebooks. Stage masks can be automatically exported as MOCs and displayed in an ipyaladin viewer layer, enabling users to pan, zoom, and overlay catalogs or images directly from the notebook. This interface is particularly useful for examining the correctness of star-exclusion radii, verifying geometry boundaries, or cross-checking against sky images.
@@ 2.7 IO and Interoperability
The Skykatana FITS format consists of a binary table with columns COVPIX (coverage pixel index), ENC (encoding flag), and PACKED (bit-packed byte array of child pixels). The table header records the sparse and coverage geometry and encoding details using the keywords listed in Table 1. These keywords enable third-party readers to interpret the bit-packed geometry unambiguously using standard FITS tools. To facilitate interoperability, Skykatana supports exporting masks to other formats (e.g. MOC). Moreover, because in-memory stages are represented as HEALSparse maps, they can also be exported to the native HEALSparse FITS format when full compatibility with external HEALSparse workflows is required.
In addition to the FITS stage files, each pipeline directory contains a JSON metadata file that (i) records the parameters and methods used to generate each stage, including the number of valid pixels and total masked area, and (ii) stores the information needed to restore the pipeline state and reattach stages under their original names when reloaded.
Keyword Meaning NSIDE_COV NSIDE of the coverage grid NSIDE_SPA NSIDE of the sparse grid NFINE Children per coverage pixel: $(\mathrm{NSIDE\_SPA}/\mathrm{NSIDE\_COV})^{2}$ DTYPE Stored logical type (currently ’bool’) ENCOD Encoding scheme (currently ’BITPACK’) BITORD Bit order within bytes (currently ’L’ for little-endian)
Table 1: Skykatana stage FITS header keywords.
@@ 3.1 Data pre-processing
To calibrate the stellar exclusion radii, we used Rubin DP1 photometry (NSF-DOE Vera C. Rubin Observatory 2025b) produced by LSST pipeline (Bosch et al., 2017) within the Extended Chandra Deep Field South (ECDFS). Among the DP1 fields, ECDFS was selected because it is the deepest LSST region currently available and targets a relatively clean extragalactic area, making it representative of low-crowding conditions relevant for cosmological analyses. In such regions, source detection and completeness are less affected by overlapping stellar halos, enabling a more robust characterization of radial density profiles around bright stars. By contrast, in highly crowded fields we expect these profiles to be noisier due to the effect of overlapping halos and source confusion, which would complicate the determination of well-defined exclusion radii.
Our first goal was to construct a background distribution of sources with approximately uniform spatial completeness. For this, we made use of the deepCoadd_psf_maglim_consolidated_map_weighted_mean property maps (Bechtol et al., 2025) available in the Rubin Science Platform222dp1.lsst.io/tutorials/notebook/203/notebook-203-1.html, which store the depth at HEALPix resolution as HEALSparse maps. These property maps allow us to determine the maximum depth reached within the ECDFS for each band, and we adopted this value as the faintest usable magnitude for the DP1 sources. These limits correspond to 25.39, 27.38, 27.12, 26.53, 25.67, and 23.77 in the u, g, r, i, z, y bands, respectively. Next, to remove shallow border regions and ensure spatial uniformity, we used the build_prop_mask function of Skykatana to generate a boolean mask selecting only those pixels where the local depth exceeds a band-dependent threshold. The thresholds were 24.5, 26.5, 26.0, 25.5, 25.0, and 23.0 in u, g, r, i, z, and y, respectively. Only DP1 sources within these masks were included in the analysis.
@@
A well-known complication in wide-field survey pipelines is that very bright stars can produce numerous artificial detections due to over-deblending of their light profiles, both in the saturated core and in the extended PSF wings. These spurious sources arise when source detection and deblending algorithms interpret complex light structures as multiple discrete objects, leading to fragmented catalogs around bright stars (Melchior et al. 2018, Mullaney et al. 2021), producing over-densities in the vicinity of bright stars. The LSST blendedness_flag is raised when the pipeline is unable to compute a reliable blendedness metric for a given detection. This typically occurs under non-ideal photometric or geometric conditions such as saturated stellar cores, extremely steep flux gradients, irregular footprints, or severe contamination where the algorithm cannot robustly determine the relation between the flux of the source and that of its inferred parent source. The blendedness_flag is provided in the LSST Object table. In practice, this flag correlates strongly with spurious deblender fragments around bright stars, and selecting only objects with blendedness_flag == False effectively removes the majority of these artifacts. Although this cut eliminates some real faint galaxies projected over bright halos, its impact on the averaged profiles is modest. For this reason, we carried out two parallel analyses: one including all DP1 sources and thus preserving the halo-induced overdensities, and one excluding halo artifacts via blendedness_flag == False. Figure 2 shows two representative mean profiles. When all sources are included, the profile exhibits a spurious central excess, followed by a depletion where the star suppresses faint detections, and eventually a secondary excess associated with halo fragmentation. When halo artifacts are removed, the profile increases smoothly and monotonically toward unity. These qualitative differences motivate the distinct definitions of stellar radius described below.
Figure 3: Fitted radius–magnitude relations for all LSST bands in the halo and no-halo cases. For the halo case, three power-law segments are fitted per band, with boundaries at $G_{\mathrm{Gaia}}=13.6$ and $G_{\mathrm{Gaia}}=15.0$, capturing the transition from strong halo contamination to the weak-halo regime. The no-halo case is well described by a single power law per band. Best-fit parameters are listed in Tables 2 and 3.